OpenStudy (anonymous):

x^2+xy-2y^2/x^3+x^2y * x/X^2+3xy+2y^2 simplify it :)

OpenStudy (anonymous):

use brackets or draw to elaborate

OpenStudy (anonymous):

\[\frac{ x ^{2} + xy - 2y ^{2}}{ x ^{3} + x ^{2}y } \times \frac{ x }{ x ^{2} + 3xy + 2y ^{2}}\]

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

yes :) it's like that

OpenStudy (anonymous):

\[\frac{ x^2 + xy + 2y^2 }{ x(x^3 + 4x^2y+5xy^2 + 2xy^3) }\]

OpenStudy (anonymous):

that's what i ended up getting

OpenStudy (anonymous):

thank u for trying but i guess it must be more simplified

OpenStudy (anonymous):

well you can't factor the top, unless you go into the quadratic equation

OpenStudy (anonymous):

the bottom can be factored as follows:

OpenStudy (anonymous):

\[\frac{ x^2 + xy+2y^2 }{ x(x+y)^2(x+2y) }\]

OpenStudy (anonymous):

\[the numerator can be simplified as (x-y)(x+2y) than the question esily simplified than u for trying :D\]

OpenStudy (anonymous):

the numerator can be simplified as (x−y)(x+2y) than the question is easily simplified btw thank u soo much for ur help :D

OpenStudy (anonymous):

uhm, you're welcome, but i don't see how you can simplify the bottom without simplifying the top..

OpenStudy (anonymous):

yes the numerator can be simplified by trial and error (x-y)(x+2y)/x^2(x+y) * x/(x+y)(x+2y)

OpenStudy (anonymous):

after cancellation the final answer is (x-y)/(x+y)^2

OpenStudy (anonymous):

sorry, was looking at the wrong thing. yeah lol confusion